# On the reconstruction limits of complex networks

### v1

@@ A.3 Stochastic block model
and $n_{r}=\sum_{i=1}^{N}\delta_{\beta_{i},r}$ counts the number of nodes in block $r$ for the partition $\bm{\beta}$. Next, we choose the edge matrix hyperprior. This matrix can be seen as the adjacency matrix of the multigraph connecting the blocks together. In Ref. Peixoto (2017), a hierarchical SBM was used as a prior for the edge matrix where each level came with its own node partition and edge matrix that, in turn, can also be modeled by a SBM, and so on until only one block remains. Here, we focus on the simpler version of this scheme, where the edge matrix prior is simply given by a multigraph ER model with $b$ nodes:
@@
Appendix B Markov chain likelihoods
@@
$P(\bm{X}_{t+1}|\bm{X}_{t},G=g)=\prod_{i=1}^{N}P(X_{i,t+1}|X_{i,t},\{X_{j,t}:j\in\mathcal{N}_{i}\})\,,$ (46)
where $\mathcal{N}_{i}$ contains the neighbors of node $i$ in the graph $g$. For time homogeneous Markov chains, the transition probability of a node $i$ is expressed in terms of its number of active neighbors $n_{i,t}$, inactive neighbors $m_{i,t}$ and some set of parameters $\varphi$. We denote the activation and deactivation probabilities $\alpha(n_{i,t},m_{i,t},\varphi)$ and $\beta(n_{i,t},m_{i,t},\varphi)$, respectively. Putting everything together, the transition probability of the Markov chain is given by
@@
In general, we fix $\alpha_{0}=\beta_{0}=0$ for the synthetic experiments, and infer them in Sec. V.1.
@@
Glauber Glauber (1963) $J$ $\sigma[2J(n-m)]$ $\sigma[2J(m-n)]$
@@
Cowan Cowan (1990) $(a,\beta,\mu,\nu)$ $\sigma[a(\nu m-\mu)]$ $\beta$
@@
The parameters $\phi$ of these models are fixed except in Sec. V.1 where they are inferred by sampling from the joint posterior $P(G,\phi,\theta|X)$. In these experiments, we use non-informative uniform prior densities for all parameters in $\phi$, and we constrain their value in finite intervals. For probability parameters, such as $\alpha_{0}$ and $\beta$ for the SIS and Cowan models, the prior density is $\rho(\phi)=1$. For positive unbounded parameters, such as $J$ for the Glauber model, and $\mu$ and $\nu$ for the Cowan model, we set the maximum value to $10$ such that $\rho(\phi)=\frac{1}{10}$. Note that we fix $a=1$ in the case of the Cowan model and $\beta_{0}=0$ for the SIS and the Cowan models in Sec. V.1, without loss of generality since they are redundant parameters that may lead to non-identifiability issues.
Appendix C Heuristic reconstruction algorithms
@@
Appendix D Proof of the optimality of the true data generating model
In this section, we prove that the true data generating (TDG) model $M^{*}=(G^{*},X^{*})$ is optimal. To do so, we start by adapting the standard data processing inequality Cover and Thomas (2006), and carry on with the optimality proof. First, let $M=(G,X)$ be a generative process with a graph prior $P(G)$ and likelihood $P(X|G)$, whose posterior probability is $P(G|X)$. Assuming that the data is generated from the generative process $M^{*}=(G^{*},X^{*})$, the probability to generate a graph $g^{*}$ followed by some data $x^{*}$, and to reconstruct a graph $g$ from $x^{*}$ is given $P(G^{*}=g^{*},X^{*}=x^{*},G=g|X=x^{*})$ [Eq. (2)]. This conditional probability is properly normalized on all values of $(g^{*},x^{*},g)$.
D.1 Modified data processing inequality
To adapt the data processing inequality in this contxt, we start from the mutual information $I(G^{*};X^{*},G|X=X^{*})$. Using the chain rule yields
$\begin{split}I(G^{*};X^{*},G|X=X^{*})&=I(G^{*};X^{*}|X=X^{*})+I(G;X^{*}|G^{*},X=X^{*})\\ &=I(G^{*};G|X=X^{*})+I(G^{*};X^{*}|G,X=X^{*})\,.\end{split}$ (54)
Many simplifications can be made. First, since $G$ and $X^{*}$ are conditionally independent given $G^{*}$, then $I(G;X^{*}|G^{*},X=X^{*})=0$. Second, because both $G^{*}$ and $X^{*}$ are independent from $X$, then $I(G^{*};X^{*}|X=X^{*})=I(G^{*};X^{*})$. Put together, we get the following data processing equality
$I(G^{*};G|X=X^{*})+I(G^{*};X^{*}|G,X=X^{*})=I(G^{*};X^{*})\,.$ (55)
Also, since $I(G^{*};X^{*}|G,X=X^{*})\geq 0$, we get the modified data processing inequality:
$I(G;\hat{G}|\hat{X}=X)\leq I(G;X)\,,$ (56)
where
$\begin{split}I(G^{*};G|X=X^{*})&=H(G^{*}|X=X^{*})-H(G^{*}|G,X=X^{*})\\ &=H(G^{*})-H(G^{*}|G,X=X^{*})\end{split}$ (57)
measure the performance of $G$ at reconstructing $G^{*}$ and $I(G^{*};X^{*})=H(G^{*})-H(G^{*}|X^{*})$ is the upper bound [Eq. (16)].
D.2 Optimality of the true data generating model
The maximum value of $I(G^{*};G|X=X^{*})$ is reached when the reconstruction model $M$ is the TDG, i.e., $M=M^{*}$, where $I(G;\hat{G}|\hat{X}=X)=I(G;X)$. To show this, we start from Eq. (55), where the second term of the RHS is again obtained from the chain rule:
$I(G^{*};X^{*}|G,X=X^{*})=H(X^{*}|G,X=X^{*})-H(X^{*}|G^{*})\,$ (58)
where the second term of the RHS was readily simplified to $H(X^{*}|G^{*},G,X=X^{*})=H(X^{*}|G^{*})$. Also, the difference is null when $H(X^{*}|G,X=X^{*})=H(X^{*}|G^{*})$, which happens when $M$ is equal in distribution to $M^{*}$. Therefore, the TDG model is optimal. Note also that this statement is not bidirectional: there may be other models that are optimal as well.
Appendix E Relationship between the posterior loss and the reconstructability
In this section we show that the posterior loss is related to the reconstructability, under certain conditions. Let $(g^{*},x^{*})$ be generated by the TDG model $M^{*}=(G^{*},X^{*})$, and $M=(G,X)$ be the reconstruction model. We define $p_{i}(x)=P(G=g_{i}|X=x)$ be the posterior probability of the graph $g_{i}$ given some observation $x$. We also denote $\bm{p}(x)=\big{(}p_{1}(x),p_{2}(x),...,p_{|\mathcal{G}|}(x)\big{)}$ the vector of the posterior probabilities of all graphs in $\mathcal{G}$. The posterior loss $L\big{(}\bm{y},\bm{p}(x)\big{)}$ measures the accuracy of the posterior probabilities $\bm{p}(x)$ at predicting the correct labels $\bm{y}$, where $y_{i}=\delta_{g^{*},g_{i}}$ is a one-hot encoding of the true graph $g^{*}$ using a Kronecker delta. It is defined as
$L\big{(}\bm{y},\bm{p}\big{)}=-\sum_{i=1}^{|\mathcal{G}|}y_{i}\log p_{i}(x)\,,$ (59)
@@ Appendix E Relationship between the posterior loss and the reconstructability
When the posterior probability factors with respect to the edges and the graphs do not contain multi-edges, i.e.,
$P(G=\bm{a}|X=x)=\prod_{i<j}\pi_{ij}(x)^{a_{ij}}\Big{(}1-\pi_{ij}(x)\Big{)}^{1-a_{ij}}\,$ (61)
the posterior loss is equal to the Eq. (20).
@@
Hence, when $M$ and $M^{*}$ are equal in distribution, $\mathbb{E}_{X^{*},G^{*}}\quantity[L(\bm{y},\bm{p}^{*})]=H(G^{*}|X^{*})$, which is the posterior entropy of the TDG model. Furthermore, the expected posterior loss is linearly related to the reconstructability, with a proportionality factor given by the negative entropy of $G$
@@
Appendix F Bounds of the information gain
In this section, we show that the information gain is non-negative and bounded by the CE between the posterior and the prior. Recall that the information gain is given by Eq. (IV.1) (equivalently Eq. (28)):
@@ Appendix F Bounds of the information gain
Appendix G Numerical approximations of the mutual information
@@
G.1 Graph enumeration approach
@@
G.2 Variational mean-field approximation
@@
where $\pi_{ij}(x)\equiv P(A_{ij}=1|X=x)$ is the marginal conditional probability of existence of the edge $(i,j)$ given $x$. For multigraphs, we obtain a similar expression involving a probability $\pi_{ij}(m\mid x)=P(A_{ij}=m|X=x)$ that there are $m$ multi-edges between $i$ and $j$. In this case, the MF posterior becomes
@@
G.3 Graph evidence estimation for the stochastic block model
@@
G.4 Evidence estimation for model selection
@@
$\begin{split}\log P(X=x)\simeq&\frac{1}{M}\sum_{k=1}^{M}\bigg{[}\log P(G=g^{(k)}|X=x)\\ &-H(G|X=x)\bigg{]}\,.\end{split}$ (81)
@@
$\begin{split}\log P(X=x)\gtrsim&\frac{1}{M}\sum_{m=1}^{M}\bigg{[}\log P(G=g^{(k)}|X=x)\\ &-\sum_{i<i}h\big{(}\pi_{ij}(x)\Big{)}\bigg{]}\,\end{split}$
@@
where $H(\phi|X)=H(\phi|G,\theta,X)$ and $H(\theta|G,X)=H(\theta|G)$, by virtue of the facts that $\phi$ is conditionally independent of $G$ and $\theta$, and that $\theta$ is conditionally independent of $X$. For evaluating $H(\phi|X)$, since $\phi$ are continuous random variables, we estimate the posterior density with kernel density estimation (KDE) with a Gaussian kernel and estimate the differential entropy from the estimated density. In the case of $H(\theta|G)$, this term only concerns the SBM prior in our experiments, where $\theta$ are discrete variables $\bm{b}$ where $b_{i}$ denotes the membership of node $i$ to a group. We use the procedure described in Sec. G.3 to estimate $H(\bm{b}|G)$.
Appendix H Markov chain Monte-Carlo algorithm
@@
This sampling scheme can be generalized when additional hyperparameters $\theta$ of the graph prior or parameters $\phi$ from the likelihood must be inferred as in Sec. V. We consider a Gibbs sampling scheme where each random variable $G$, $\theta$ and $\phi$ is sampled sequentially and conditioned on the others. In all cases, the acceptance probability follows Eq. (84), where $G$ is replaced by either $\theta$ or $\phi$ when these parameters are sampled. In this paper, only the SBM among the considered graph models contains parameters of the type of $\theta$. To sample from these, we use the same procedure as in Ref. (Peixoto, 2017, Sec. VI)—we refer to it for further detail. The data models considered contains many parameters that we would like to infer, for example, the infection and recovery probabilities, $\lambda$ and $\beta$, respectively, in the SIS. These parameters are real number constrained within an interval (for instance, $[0,1]$ for the recovery probability $\beta$); Hence, any proposed move where $\varphi$ falls outside of this interval is rejected. We propose moves drawn from a Gaussian distribution with density
@@
In Sec. V, we fix $\sigma=0.1$.
Appendix I Reconstructability of graph models with delta distribution
Suppose $X$ is generated using a single graph $g^{*}$. If we were to observe many realizations of $X$ with the single graph $g^{*}$, the graph prior of the TDG would be $P(G=g)=\delta_{g,g^{*}}$ and the evidence of this process would be exactly equal to the likelihood of the TDG process, denoted $p^{*}(X)\equiv P(X|G=g^{*})$. As a result, the mutual information and the entropy of $G$ would both be zero, and so the reconstructability would be undefined.
@@ Appendix I Reconstructability of graph models with delta distribution
Appendix J Inference of brain networks
Fig. 8: Posterior of the maximum evidence model (SIS model with SBM prior): (a) posterior probability matrix of the edge occupancy, (b) histogram of the infection probability, (c) the recovery probability and (d) the auto-activation probability. In (a), each entry of the matrix represents the number of times the edge has been sampled, among the 1000 posterior samples. Also, we highlight the probe partition of the graph using deemed black separation lines.
@@
In this section, we describe the procedure we used to reconstruction the mouse brain network from Sec. V.1. The raw data is available in Steinmetz et al. (2019), which was originally presented in Ref. Stringer et al. (2019). We refer to their paper for any technical detail regarding the data collection.
J.1 Data preprocessing
@@
J.2 Inference procedure
The inference procedure is very similar to that presented in Appendix H. We consider a model parametrized by $\hat{X}$, $\hat{G}$ and their parameters $\hat{\phi}$ and $\hat{\theta}$, respectively. However, we have an additional limitation: We do not know the number of edges in the graph. We tried extending our MCMC algorithm by including moves that do not preserve the number of edges—i.e., adding or removing a single edge—, but we found that these attempts suffered from poor mixing time.
To alleviate this problem, we propose to search for the number of edges first, by minimizing the description length $\log P(\hat{X},\hat{\phi},\hat{G},\hat{\theta})$ Peixoto (2024). We solve this optimization problem using a semi-greedy algorithm where we propose $K$ move candidates, and select the one that locally minimizes the objective. Like for our MCMC algorithm, we iterate over $\hat{G}$, $\hat{\theta}$ and $\hat{\phi}$ sequentially to locally perform the optimization on each of them independently. At each step, we sample $10000$ candidates for $\hat{G}$ and $\hat{\theta}$, and $10$ for $\hat{\phi}$. Once the number of edges has converged, we stop the semi-greedy algorithm and freeze the number of edges. The MCMC algorithm then proceeds to sample from the posterior with a fixed number of edges.
J.3 Posterior inspection

### v2

@@ IV.1.3 Stochastic block model
and $n_{r}=\sum_{i=1}^{N}\delta(\beta_{i},r)$ counts the number of nodes in block $r$ for the partition $\bm{\beta}$. Next, we choose the edge matrix hyperprior. This matrix can be seen as the adjacency matrix of the multigraph connecting the blocks together. In Ref. Peixoto (2017), a hierarchical SBM was used as a prior for the edge matrix where each level came with its own node partition and edge matrix that, in turn, can also be modeled by a SBM, and so on until only one block remains. Here, we focus on the simpler version of this scheme, where the edge matrix prior is simply given by a multigraph ER model with $b$ nodes:
@@
IV.2 Markov chain likelihoods
@@
$P(\bm{X}_{t+1}|\bm{X}_{t},G=g)=\prod_{i=1}^{N}P(X_{i,t+1}|X_{i,t},X_{\mathcal{N}_{i},t})\,,$ (47)
where $X_{\mathcal{N}_{i},t}\equiv(X_{j,t})_{j\in\mathcal{N}_{i}}$ contains the state of the neighbors $\mathcal{N}_{i}$ of node $i$ in the graph $g$. For time homogeneous Markov chains, the transition probability of a node $i$ is expressed in terms of its number of active neighbors $n_{i,t}$, inactive neighbors $m_{i,t}$ and some set of parameters $\varphi$. We denote the activation and deactivation probabilities $\alpha(n_{i,t},m_{i,t},\varphi)$ and $\beta(n_{i,t},m_{i,t},\varphi)$, respectively. Putting everything together, the transition probability of the Markov chain is given by
@@
In general, we fix $\alpha_{0}=\beta_{0}=0$ for the synthetic experiments, and infer them in Sec. II.4.1.
@@
Glauber Glauber (1963) $J$ $\sigma\quantity(2J(n-m))$ $\sigma\quantity(2J(m-n))$
@@
Cowan Cowan (1990) $(a,\beta,\mu,\nu)$ $\sigma\quantity(a(\nu m-\mu))$ $\beta$
@@
The parameters $\phi$ of these models are fixed except in Sec. II.4.1 where they are inferred by sampling from the joint posterior $P(G,\phi,\theta|X)$. In these experiments, we use non-informative uniform prior densities for all parameters in $\phi$, and we constrain their value in finite intervals. For probability parameters, such as $\alpha_{0}$ and $\beta$ for the SIS and Cowan models, the prior density is $\rho(\phi)=1$. For positive unbounded parameters, such as $J$ for the Glauber model, and $\mu$ and $\nu$ for the Cowan model, we set the maximum value to $10$ such that $\rho(\phi)=\frac{1}{10}$. Note that we fix $a=1$ in the case of the Cowan model and $\beta_{0}=0$ for the SIS and the Cowan models in Sec. II.4.1, without loss of generality since they are redundant parameters that may lead to non-identifiability issues.
IV.3 Heuristic reconstruction algorithms
@@
IV.4 Relationship between the posterior loss and the reconstructability
In this section we show that the posterior loss is related to the reconstructability, under certain conditions. Let $(g^{*},x^{*})$ be generated by the TDG model $M^{*}=(G^{*},X^{*})$, and $M=(G,X)$ be the reconstruction model. We define $p_{i}(x)=P(G=g_{i}|X=x)$ be the posterior probability of the graph $g_{i}$ given some observation $x$. We also denote $\bm{p}(x)=\big{(}p_{1}(x),p_{2}(x),...,p_{|\mathcal{G}|}(x)\big{)}$ the vector of the posterior probabilities of all graphs in $\mathcal{G}$. The posterior loss $L\big{(}\bm{y},\bm{p}(x)\big{)}$ measures the accuracy of the posterior probabilities $\bm{p}(x)$ at predicting the correct labels $\bm{y}$, where $y_{i}=\delta(g^{*},g_{i})$ is a one-hot encoding of the true graph $g^{*}$ using a Kronecker delta. It is defined as
$L\big{(}\bm{y},\bm{p}\big{)}=-\sum_{i=1}^{|\mathcal{G}|}y_{i}\log p_{i}(x)\,.$ (55)
@@ IV.4 Relationship between the posterior loss and the reconstructability
When the posterior probability factors with respect to the edges and the graphs do not contain multiedges, i.e.,
$P(G=\bm{a}|X=x)=\prod_{i<j}\pi_{ij}(x)^{a_{ij}}\Big{(}1-\pi_{ij}(x)\Big{)}^{1-a_{ij}}\,,$ (57)
the posterior loss is given by Eq. (21).
@@
Hence, when $M$ and $M^{*}$ are equal in distribution, $\mathbb{E}_{X^{*},G^{*}}\quantity[L(\bm{y},\bm{p}^{*})]=H(G^{*}|X^{*})$, which is the posterior entropy of the TDG model. Furthermore, the expected posterior loss is linearly related to the reconstructability, with a proportionality factor given by the entropy of $G^{*}$:
@@
IV.5 Bounds of the information gain
In this section, we show that the information gain is non-negative and bounded by the CE between the posterior and the prior. Recall that the information gain is given by Eq. (II.3.1) (equivalently Eq. (30)):
@@ IV.5 Bounds of the information gain
IV.6 Numerical approximations of the mutual information
@@
IV.6.1 Graph enumeration approach
@@
IV.6.2 Variational mean-field approximation
@@
where $\pi_{ij}(x)\equiv P(A_{ij}=1|X=x)$ is the marginal conditional probability of existence of the edge $(i,j)$ given $x$. For multigraphs, we obtain a similar expression involving a probability $\pi_{ij}(m\mid x)=P(A_{ij}=m|X=x)$ that there are $m$ multiedges between $i$ and $j$. In this case, the MF posterior becomes
@@
IV.6.3 Graph evidence estimation for the stochastic block model
@@
IV.6.4 Evidence estimation for model selection
@@
$\begin{split}\log P(X=x)\simeq&\frac{1}{M}\sum_{k=1}^{M}\bigg{[}\log P(G=g^{(k)},X=x)\\ &-H(G|X=x)\bigg{]}\,.\end{split}$ (77)
@@
$\begin{split}\log P(X=x)\gtrsim&\frac{1}{M}\sum_{m=1}^{M}\bigg{[}\log P(G=g^{(k)},X=x)\\ &-\sum_{i<i}h\big{(}\pi_{ij}(x)\Big{)}\bigg{]}\,\end{split}$
@@
where $H(\phi|X)=H(\phi|G,\theta,X)$ and $H(\theta|G,X)=H(\theta|G)$, by virtue of the facts that $\phi$ is conditionally independent of $G$ and $\theta$, and that $\theta$ is conditionally independent of $X$. For evaluating $H(\phi|X)$, since $\phi$ are continuous random variables, we estimate the posterior density with kernel density estimation (KDE) with a Gaussian kernel and estimate the differential entropy from the estimated density. In the case of $H(\theta|G)$, this term only concerns the SBM prior in our experiments, where $\theta$ are discrete variables $\bm{b}$ where $b_{i}$ denotes the membership of node $i$ to a group. We use the procedure described in Sec. IV.6.3 to estimate $H(\bm{b}|G)$.
IV.7 Markov chain Monte-Carlo algorithm
@@
This sampling scheme can be generalized when additional hyperparameters $\theta$ of the graph prior or parameters $\phi$ from the likelihood must be inferred as in Sec. II.4. We consider a Gibbs sampling scheme where each random variable $G$, $\theta$ and $\phi$ is sampled sequentially and conditioned on the others. In all cases, the acceptance probability follows Eq. (80), where $G$ is replaced by either $\theta$ or $\phi$ when these parameters are sampled. In this paper, only the SBM among the considered graph models contains parameters of the type of $\theta$. To sample from these, we use the same procedure as in Ref. (Peixoto, 2017, Sec. VI)—we refer to it for further detail. The data models considered contains many parameters that we would like to infer, for example, the infection and recovery probabilities, $\lambda$ and $\beta$, respectively, in the SIS. These parameters are real number constrained within an interval (for instance, $[0,1]$ for the recovery probability $\beta$); Hence, any proposed move where $\varphi$ falls outside of this interval is rejected. We propose moves drawn from a Gaussian distribution with density
@@
In Sec. II.4, we fix $\sigma=0.1$.
IV.8 Reconstructability of graph models with delta distribution
Suppose $X$ is generated using a single graph $g^{*}$. If we were to observe many realizations of $X$ with the single graph $g^{*}$, the graph prior of the TDG would be $P(G=g)=\delta(g,g^{*})$ and the evidence of this process would be exactly equal to the likelihood of the TDG process, denoted $p^{*}(X)\equiv P(X|G=g^{*})$. As a result, the mutual information and the entropy of $G$ would both be zero, and so the reconstructability would be undefined.
@@ IV.8 Reconstructability of graph models with delta distribution
IV.9 Inference of brain networks
Fig. 8: Posterior of the maximum evidence model (SIS model with SBM prior): (a) posterior probability matrix of the edge occupancy, (b) histogram of the infection probability, (c) the recovery probability and (d) the auto-activation probability. In (a), each entry of the matrix represents the number of times the edge has been sampled, among the 8000 posterior samples. Also, we highlight the probe partition of the graph using deemed black separation lines.
@@ IV.9 Inference of brain networks
In this section, we describe the procedure we used to reconstruction the mouse brain network from Sec. II.4.1. The raw data is available in Steinmetz et al. (2019), which was originally presented in Ref. Stringer et al. (2019). We refer to their paper for any technical detail regarding the data collection.
Average Std. Dev.
Model Graph prior
Cowan ER 1 296.75 77.00
UCM 1 302.75 70.38
SBM 1 462.00 126.03
Glauber ER 18 098.88 53.94
UCM 18 612.38 93.50
SBM 18 895.63 79.47
SIS ER 1 388.50 28.92
UCM 1 296.63 50.15
SBM 1 722.38 89.98
Table 3: Statistics for the number of edges determined from the semi-greedy algorithm for each reconstruction model considered in Sec. II.4.1. The highlighted row (SIS with SBM) corresponds to the maximum evidence model associated with Figs. 8 and 9. The average and standard deviations (std. dev.) are obtained from the 8 parallel chains used for the inference.
IV.9.1 Data preprocessing
@@
IV.9.2 Inference procedure
The inference procedure is very similar to that presented in Appendix IV.7. We consider a model parametrized by $X$, $G$ and their parameters $\phi$ and $\theta$, respectively. However, we have an additional limitation: We do not know the number of edges in the graph. We tried extending our MCMC algorithm by including moves that do not preserve the number of edges—i.e., adding or removing a single edge—, but we found that these attempts suffered from poor mixing time.
To alleviate this problem, we propose to search for the number of edges first, by minimizing the description length $\log P(X,\phi,G,\theta)$ Peixoto (2024). We solve this optimization problem using a semi-greedy algorithm where we propose $K$ move candidates, and select the one that locally minimizes the objective. Like for our MCMC algorithm, we iterate over $G$, $\theta$ and $\phi$ sequentially to locally perform the optimization on each of them independently. At each step, we sample $10000$ candidates for $G$ and $\theta$, and $10$ for $\phi$. Once the number of edges has converged, we stop the semi-greedy algorithm and freeze the number of edges. The MCMC algorithm then proceeds to sample from the posterior with a fixed number of edges.
In Table 3, we summarize the results of the semi-greedy search for the number of edges. Given that the number of nodes is $1462$, our results show that the inferred networks are surprisingly sparse, except for the Glauber model which inferred one order of magnitude more edges than the Cowan and SIS models.
IV.9.3 Posterior inspection